OpenStudy (anonymous):

Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive? Numerical Answers Expected!

OpenStudy (anonymous):

Maya received 93,750 votes.

OpenStudy (anonymous):

no Numerical Answers Expected!

OpenStudy (anonymous):

you said numerical answers expected before

OpenStudy (anonymous):

i kno but its an worrd problem you have to write it out in numerical...

OpenStudy (anonymous):

Since each person can vote only once, 125,000 people means 125,000 votes, i.e. y = 125,000. x - 62,500 = 0.25x125,000 x = 31250 + 62500 x = 93,750